# Enhancing Trust in Language Model-Based Code Optimization through RLHF: A Research Design

### v1

@@ I Introduction
Performance is a crucial attribute of software code [1]. With the rapid advancement of AI, software engineering increasingly relies on AI-driven approaches, particularly language models (LMs), to enhance code performance [2, 3, 4]. However, a recent survey [5] on LM-based code optimization highlights a significant challenge: the trustworthiness and reliability of LMs. Specifically, LMs can produce hallucinations—unreliable or incorrect responses [6, 7, 8], emphasizing the need for human expertise to guide their inference, validate outputs, and ensure reliability.
To address this gap, this research aims to develop reliable, LM-powered methods for code optimization that effectively integrate human feedback. The primary goals are:
@@
    Explore different strategies for incorporating human feedback into LM-based code optimization.
@@
    Design a novel code optimization method that enhances existing approaches by leveraging human feedback.
@@ II Research Design
This project adopts a structured methodology inspired by the Research Design Canvas [9], addressing key elements outlined in Figure 1.
@@ II-B Research Topic
The research focuses on the intersection of generative AI, code optimization, and software engineering, with the goal of developing a framework that enhances the trustworthiness of LM-based code optimization with knowledge from human experts. One of the most promising research topics therein is reinforcement learning from human feedback (RLHF) [12]. Unlike traditional fine-tuning, which uses static labeled datasets to adjust model parameters, RLHF dynamically incorporates human feedback to train a reward model. This reward model guides the language model through reinforcement learning, enabling it to adapt and optimize code based on real-time human preferences and standards. Thereby, this approach enhances the capabilities of LMs in optimizing code while ensuring that human expertise and feedback play a crucial role in the process.
@@ II-F3 Access
Access to the research paper will be provided through open-source platforms to encourage widespread adoption and collaboration. For proprietary components, such as patents, access may follow the company policy to sustain the project’s impact while balancing intellectual property considerations.
@@ III Questions and Feedback Request
Finally, it would be appreciated to discuss the ethical considerations—are there any other ethical considerations in using RLHF for code optimization, and how can they be addressed? Addressing these ethical concerns could be crucial for fostering trust and acceptance of AI-driven code optimization within the developer community.
@@ IV Conclusion
This research design addresses the critical challenge of trustworthiness and reliability in LM-based code optimization by integrating human feedback into the optimization process. By exploring various strategies for incorporating human feedback and designing a novel code optimization method, this work will enhance the capabilities of language models while ensuring that human expertise plays a crucial role. Questions to discuss at DECS include methods for validating the optimized codes, mitigating biases introduced by developer preferences, and addressing ethical considerations to foster trust and acceptance within the developer community.

### v2

@@ Abstract
publicationid: pubid: Submission to the DECS of CHASE 2025. © with the authors.
@@ I Introduction
Performance is a crucial attribute of software code [1]. With the rapid advancement of AI, software engineering increasingly relies on AI-driven approaches, particularly language models (LMs), to enhance code performance [2, 3, 4]. However, a recent survey [5] on LM-based code optimization highlights a significant challenge: LMs can produce hallucinations—unreliable or incorrect responses [6, 7, 8]—which undermines the trustworthiness and reliability of LMs.
To address this gap, this research aims to develop reliable, LM-powered methods for code performance optimization that effectively integrate human feedback. The primary goals are:
@@
    Explore various strategies for incorporating human feedback into LM-based code intelligence techniques.
@@
    Design a novel code performance optimization method that leverages human feedback to enhance code performance while fostering trust in the optimization process.
@@ II Research Design
This project adopts a structured methodology inspired by the Research Design Canvas [9], as outlined in Figure 1.
@@ II-B Research Topic
The research focuses on the intersection of generative AI, code optimization, and software engineering, with the goal of developing a framework that enhances the trustworthiness of LM-based code optimization with knowledge from human experts. A promising research topic therein is reinforcement learning from human feedback (RLHF) [12]. Unlike traditional fine-tuning, which uses static labeled datasets to adjust model parameters, RLHF dynamically incorporates human feedback to train a reward model. This reward model guides the language model through reinforcement learning, enabling it to adapt and optimize code performance based on real-time human preferences and standards. Thereby, this approach enhances the capabilities of LMs in optimizing code performance while ensuring that human expertise and feedback play a crucial role in the process.
@@ II-F3 Access
Access to the research paper will be provided through open-source platforms to encourage open science. For proprietary components such as patents, access may follow the company policy to sustain the project’s impact while balancing intellectual property considerations.
@@ III Questions and Feedback Request
Finally, given the multiple approaches to enhancing cooperative and human-centric aspects in LM-based code performance optimization tools—such as using developer feedback as input, improving developer-tool interaction, and assessing the tool’s impact on coding practices—what are additional strategies that could more effectively build trust in LM-based code optimization? Insights on how to balance these methods or adopt alternative approaches would be invaluable.
@@ IV Conclusion
This research design addresses the critical challenge of trustworthiness and reliability in LM-based code performance optimization by integrating human feedback into the optimization process. By exploring various strategies for incorporating human feedback and designing a novel code optimization method, this work will enhance the performance optimization capabilities of language models while ensuring that human expertise plays a crucial role. Questions to discuss at DECS include methods for validating the optimized codes, mitigating biases introduced by developer preferences, and identifying alternative approaches to foster trust and acceptance within the developer community.